OpenStudy (anonymous):

how do i reduce[ (x)(x)^4]

OpenStudy (anonymous):

Using the properties of exponents, you can see straightaway that \[x*x^4=x*(x*x*x*x)=x*x*x*x*x=x^5\]

OpenStudy (anonymous):

x = x^1